OpenStudy (anonymous):

Write the equation of the line that satisfies the given conditions. Express the final equation in standard form. Contains the point (-8, 7) and is parallel to the line x − 7y = 6

OpenStudy (whpalmer4):

Here are the steps you need to take: 1) find the slope of the line \(x -7y=6\) 2) use the point-slope formula to write the equation of a new line which contains the point \((-8,7)\) and the slope we found in step 1. 3) rearrange the equation as necessary to put it in standard form \(Ax+By = C\) To find the slope of a line, put it in slope-intercept form \(y = mx + b\) and read off the slope \(m\) The point-slope formula is \[y-y_0 = m(x-x_0)\]where \((x_0,y_0)\) is the known point